A fox locates rodents under the snow by the slight sounds they make. The fox then leaps straight into the air and burrows its nose into the snow to catch its meal. If a fox jumps up to a height of 91 cm , calculate the speed at which the fox leaves the snow and the amount of time the fox is in the air. Ignore air resistance.
Answer #1

a)

Using 3rd equation of motion

v^2= u^2+ 2ad

v^2= 2* 9.8*0.91

v= 4.22 m/s

==========

v= u + gt

t= 4.22/9.8

t= 0.4306 second

Total time= 2*0.43= 0.86 seconds